Jacob bought a scooter for a certain sum of money. He spent 10% of the cost, on repairs and sold the scooter for a profit of RS 1100. How much did he spend on repairs if he made a profit of 20%

Grade:

1 Answers

Ravi
9 years ago
Let Jacob have bought the scooter at x
Repairing cost= 10% of x= x/10

Net CP= x+x/10=1.1x

So 20% of 1.1x=1100

1.1x(20/100)=1100
Find x.

Find the cost of repair(i.e. 0.1 x)
..................................................................................................................................